# CLIFFORD ALGEBRAS, MESON ALGEBRAS AND HIGHER ORDER GENERALIZATIONS

### v1

@@ 1 Introduction
where $T(E)$ is the tensor algebra of $E$ and where $(I)$ for $I\subset T(E)$ denotes the 2-sided ideal of $T(E)$ generated by $I$.
@@ 2 The neutral meson algebra $D_{0}(E)$
$D^{n}_{0}(E)=\left\{\begin{array}[]{ll}D^{2m}_{0}(E)=\wedge^{m}E\otimes\wedge^{m}E&\mbox{for}\>\>n=2m\\ D^{2m+1}_{0}(E)=\wedge^{m+1}(E)\otimes\wedge^{m}E&\mbox{for}\>\>n=2m+1\end{array}\right$. (5)
@@
The isomorphism classes of irreducible representations of $GL(E)$ are labeled by the Young diagrams $\lambda$ with at most $N$ lines (i.e. with columns of length $\leq N$).
@@
$\begin{array}[]{lll}\sum_{k\geq 0}{\mbox{dim}}(\lambda_{m+k,m-k})&=(C^{m}_{N})^{2}-C^{m+1}_{N}C^{m-1}_{N}\\ \\ &+(C^{m+1}_{N}C^{m-1}_{N}-C^{m+2}_{N}-C^{m-2}_{N})+\cdots\\ \\ &+(C^{m+k}_{N}C^{m-k}_{N}-C^{m+k+1}_{N}C^{m-k-1})+\cdots\end{array}$
@@ 3.2 The remarkable algebra ${\mathcal{F}}(E)$
where $\lambda$ runs over all symmetric diagrams with columns of length smaller or equal to $N$, ($\lambda\mapsto\bar{\lambda}$ denotes the symmetry with respect to the main diagonal).

### v2

@@ 1 Introduction
where $T(E)$ is the tensor algebra of $E$ and where $(I)$ for $I\subset T(E)$ denotes the 2-sided ideal of $T(E)$ generated by II.
@@ 2 The neutral meson algebra $D_{0}(E)$
$D^{n}_{0}(E)=\left\{\begin{array}[]{ll}D^{2m}_{0}(E)=\wedge^{m}E\otimes\wedge^{m}E&\mbox{for}\>\>n=2m\\ D^{2m+1}_{0}(E)=\wedge^{m+1}E\otimes\wedge^{m}E&\mbox{for}\>\>n=2m+1\end{array}\right$. (5)
@@
The isomorphism classes of irreducible representations of $GL(E)$ are labeled by the Young diagrams $\lambda$ with at most NN lines (i.e. with columns of length $\leq N$).
@@
$\begin{array}[]{lll}\sum_{k\geq 0}{\mbox{dim}}(\lambda_{m+k,m-k})&=(C^{m}_{N})^{2}-C^{m+1}_{N}C^{m-1}_{N}\\ \\ &+(C^{m+1}_{N}C^{m-1}_{N}-C^{m+2}_{N}C^{m-2}_{N})+\cdots\\ \\ &+(C^{m+k}_{N}C^{m-k}_{N}-C^{m+k+1}_{N}C^{m-k-1})+\cdots\end{array}$
@@ 3.2 The remarkable algebra ${\mathcal{F}}(E)$
where $\lambda$ runs over all symmetric diagrams with columns of length smaller or equal to NN, ($\lambda\mapsto\bar{\lambda}$ denotes the symmetry with respect to the main diagonal).